OpenStudy (anonymous):

A 75.0 kg person steps on a scale in an elevator. The scale reads 761 N. What is the magnitude of the acceleration of the elevator?

OpenStudy (stormfire1):

The scale will read the total forces acting down on it (the sum of the normal forces). The forces pushing up on the person will be mg + m(elevator's acceleration) which will be opposed by a normal force of the same magnitude. So the equation would be: \[Ftotal=761N=mg+Fn=(75.0kg∗9.8m/s2)+(75kg∗a)\]Now just solve for a.

OpenStudy (stormfire1):

Also, another way to look at it is: \[\frac{761N}{75.0kg}= 10.1m/s^2\]That would be the elevator's acceleration if there were no gravitational forces acting on it. So subtract the force of gravity from that to get:\[10.1m/s^2-9.8m/s^2 = 0.34 m/s^2 (upward)\]Either way you get the same answer.

OpenStudy (anonymous):

the computer is not accepting it as my answer

OpenStudy (stormfire1):

I really can't help you there...I'm not familiar with how you're supposed to input it. It may accept 10.1 m/s^2 but that's really not correct. It should be either 0.3 m/s^2 or 0.34 m/s^2...depending on how you do the calculations

OpenStudy (stormfire1):

I rounded my calcs based on significant figures...you could try doing them again without any rounding

OpenStudy (anonymous):

ok it accepted 0.333m/s^2 as the answer,i took g=9.81

OpenStudy (anonymous):

but why shuld we subtract from 9.81? cos if i were to take this in exam,i won't knw am to do dat

OpenStudy (stormfire1):

You have to subtract 9.81 because the force of gravity is also pulling down on the scale